Assume the following model of the expenditure sector:

C = 420 + (4/5)YD     

T0 = 100                            

Io = 160                       

Go = 180                      

NXo = - 40

1- Calculate the multiplier?

2- Calculate the equilibrium GDP?

3- If the government would like to increase the equilibrium level of output (Y) to the full-employment level Y* = 3,800, by how much should government purchases (G) be changed?

4- Suppose government imposes a progressive tax policy and the new tax represents by:

T = 100 + 1/4 Y

a- Calculate the new multiplier?
b- Calculate the new equilibrium GDP?
Answer #1

1). 2a - . 5 1 1-1 45 1-0.2 0.2 Y = 420 + 4y4 (100) + 160 +180 yo 5 5 4/5 = 640 = y* - 3200 y = 3200 La = or = 5= 3800-3200 O4. NT 100 + 7 -0.8(1-0.25) 2.5 1-0-810.75 6.4 Y420 + ” (Y-100-) +160+180 - 40 P Y 5 4 34 = 5 4 640 7 Y-37 = 640 Y*) 2Y = 640